question 12 Part A 1.2 kg piece of ice at 0 °C is thrown into a very large What is the entropy change of the ice? lake at 10 °c. Water: Lf= 334 kukg c = 4.19 kJ/(kg Co) J/K Submit My Answers Give Up Part B What is the total entropy change of the ice and lake system? J/K Submit My Answers Give Up Provide F Feedback Continue
Answer #1

A) Since the lake is very large, it's temperature will not change. Only the ice will melt. The energy involved in this is:

Q = mL = (1.2)(334) = 400.8 kJ

Temperature of ice T = 0 degree C = (0+273) K = 273 K

Change in entropy of ice = Q/T = 400.8/273 = 1.4681 kJ/k or 1468.3 J/K (ans)

B)

Temperature of system T = 10 degree C = (10+273) K = 283 K

so, Change in entropy of the system = Q/T = 400.8/283 = 1.4163 kJ/k or 1416.3 J/K (ans)